Solve: 2x+3/9 is less than or equal to x-1/6
Answer by checkley77(12844)

2x+3/9<=x-1/6
2x+1/3<=x-1/6
2x-x<=-1/6-1/3
x<=(-1-2*1)/6
x<=(-1-2)/6
x<=-3/6
x<=-1/2 answer.
proof:
let x=-1
2*-1+3/9<=-2-1/6
-2+1/3<=(-2*6-1)/6
(-2*3+1)/3<=(-12-1)/6
(-6+1)/3<=-13/6
-7/3<=-13/6
-2.333<=2.167